An echocardiogram uses 4.4 MHz ultrasound to measure blood flow in the aorta. The blood is moving away from the probe at 1.6 m/s .

A. What is the frequency shift of the reflected ultrasound? Use values given in the table if necessary. Enter a positive value if the received frequency is higher than the emitted frequency, and a negative value if the received frequency is lower than the emitted frequency.
Answer #1

Frequency, F = (v - vblood)/(v + vblood) f

F = (1540 - 1.6)/(1540 + 1.6) x 4.4 x 10^6 = 4.39 x 10^6 Hz

Shift = (4.39 x 10^6) - (4.4 x 10^6) = - 9133.37 Hz

Comment in case any doubt please rate my answer....